The last four weekly values of sales were 80, 100, 105, and 90 units, respectively. The...

  1. The last four weekly values of sales were 80, 100, 105, and 90 units, respectively. The last four forecasts (for the same four weeks) were 60, 80, 95, and 75 units, respectively. Calculate the MAD, MSE, and MAPE for these four weeks.

Sales

Forecast

∣ Error ∣

Error squared

Pct (%). error

80

60

100

80

105

95

90

75

MAD

MSE

MAPE
Answer #1

MAD(Mean Absolute Deviation) = Sum of absolute deviations/Number of periods

= {|80-60| + |100-80| + |105-95| + |90-75|}/4

= (20+20+10+15)/4

= 16.25

MSE(Mean Square Error) = (202+202+102+152)/4

= (400+400+100+225)/4

= 281.25

MAPE(Mean Absolute Percentage Error) = Sum of absolute percentage errors/Number of periods

= [{(80-60)/80}*100 + {(100-80)/100}*100 + {(105-95)/105}*100 + {(90-75)/90}*100]/4

= (25 + 20 + 9.52 + 16.67)/4

= 17.8%